The qoutient of two numbers is 8 and their sum is 90. Find the number ?
Answer by oscargut(2103)

The qoutient of two numbers is 8 and their sum is 90. Find the number ?
let name the numbers x and y then:
1) x/y = 8 (The qoutient of two numbers is 8)
2) x+y = 90 (their sum is 90)
using 1st eq x=8y then substituting in the 2nd 8y+y=90
then 9y = 90 then y = 10 then x = 8y = 8(10) = 80
R: numbers are 80 and 10